Please I'm asking for your help to solve the problem.
Question
The records of a department store show that 10% of their customers who make a purchase return the merchandise in order to exchange it. What is the probability that in the next eight purchases at least three customers return the merchandise for exchange?

Answer by rothauserc(4718)

Use the binomial probability distribution
:
Probability(P) (k successes out of n trials) = nCk * p^k * (1-p)^(n-k), where n is the number of trials, k is the number of successes, p is probability of success, nCk is combination of n trials taken k at a time( n! / k! * (n-k)!)
:
P(in 8 purchases at least 3 customers return purchases) = 1 -P(0 purchase returned) -P(1 purchase returned) - P(2 purchases returned)
:
Note that n = 8, p = 0.10, k is 0, 1, 2
:
*******************************************************************************
P(in 8 purchases at least 3 customers return purchases) = 1 -0.43 -0.38 -0.15 = 0.04
*******************************************************************************
:
Note this solution assumes that each customer makes 1 purchase
: